JNTU

JNTU B.Tech III RELIABILITY ENGINEERING Previous Question Paper (November/December - 2018 - R16)

B.Tech | III | III-I | R16 | November/December - 2018

Answer PaperPaper Code: 135CYMaximum Marks: 75Time: 3 hours

Part A is compulsory. Answer all questions in Part A. Part B consists of 5 Units. Answer any one full question from each unit.

Part A

125 marks
a

Define the terms Availability and Reliability.

2 marks

Answer

Reliability is the probability that a component or system will perform its required function under specified conditions for a stated period of time. Availability is the probability that a system is operating satisfactorily at any given time, accounting for both its reliability and maintainability.

b

What is the probability of getting a total of 9 spots in a single throw of 2 dice?

3 marks

Answer

The total number of outcomes when throwing two dice is 6 × 6 = 36. The combinations that result in a sum of 9 are (3,6), (4,5), (5,4), and (6,3). There are 4 favorable outcomes.

P(9) = 4 / 36 = 1 / 9 ≈ 0.1111
c

A system consists of 15 identical components, all of which must work for system success. What is the system reliability if each component has a reliability of 0.9?

2 marks

Answer

Sign in to view the complete answer paper.

d

An engineering system design requires 100 similar components in series. What is the minimum reliability of each system if the overall reliability must be at least 0.95?

3 marks

Answer

Sign in to view the complete answer paper.

e

Define failure density function. Write the expression for it.

2 marks

Answer

Sign in to view the complete answer paper.

f

Write hazard rate function equation in terms of failure density function f(t) and reliability function R(t).

3 marks

Answer

Sign in to view the complete answer paper.

g

Define Stochastic Transitional Probability Matrix (STPM).

2 marks

Answer

Sign in to view the complete answer paper.

h

Write the expressions for limiting state probabilities for a continuous Markov process.

3 marks

Answer

Sign in to view the complete answer paper.

i

Define frequency of encountering a state and write its equation.

2 marks

Answer

Sign in to view the complete answer paper.

j

What is a Cutset? What is its use in the reliability evaluation of engineering systems?

3 marks

Answer

Sign in to view the complete answer paper.

Part B

210 marks
a

List and discuss different rules for combining the probabilities.

7 marks

Answer

Sign in to view the complete answer paper.

b

One card is drawn from a standard pack of 52 playing cards. Let ‘X’ be the event that it is a red card and ‘Y’ be the event that it is a court or face card. What is the probability that both ‘X’ and ‘Y’ occur?

3 marks

Answer

Sign in to view the complete answer paper.

OR
310 marks
a

Prove that the expected value and standard deviation of an exponential distribution are equal.

6 marks

Answer

Sign in to view the complete answer paper.

b

Define and explain the terms: Mean Time to Failure, Mean time to Repair and Mean Time between Failures.

4 marks

Answer

Sign in to view the complete answer paper.

4

Derive a general expression for the unreliability of the system whose reliability model is shown in figure 1. Consider the case in which all parallel branches of this system are fully redundant with an exception of that consisting of components D, E and F for which any 2 of the branches are required for system success. Also find the system reliability if all components have the same reliability of 0.9.

10 marks

Answer

Sign in to view the complete answer paper.

OR
5

The success of the system shown in the below figure 2 requires that at least one of the paths, AC, BD, AED, BEC is good. Evaluate a general expression for system success and the reliability of the system if each component has a reliability of 0.85.

10 marks

Answer

Sign in to view the complete answer paper.

610 marks
a

Define hazard rate.

2 marks

Answer

Sign in to view the complete answer paper.

b

The hazard rate of a device is λ(t) = 1/t. Deduce i) The probability density function ii) The survivor function iii) The expected value, and iv) The variance.

8 marks

Answer

Sign in to view the complete answer paper.

OR
7

A system is composed of three elements. Component 1 is in series with the parallel redundant combination of components 2 and 3, which are identical. λ1 = 1 f/yr, λ2 = λ3 = 3 f/yr and µ1 = 98 r/yr, µ2 = µ3 = 50 r/yr.

10 marks
a

Calculate the reliability of this system for a mission of 1500 hr.

5 marks

Answer

Sign in to view the complete answer paper.

b

What is the mean time to failure for this system?

5 marks

Answer

Sign in to view the complete answer paper.

8

The state space diagram and transitional rates in f/hr of a continuous Markov process is shown in figure 3. Calculate: a) The limiting probabilities of each state b) The availability of the system.

10 marks

Answer

Sign in to view the complete answer paper.

OR
910 marks
a

Explain what is meant by limiting state probability and how it is evaluated for discrete Markov chain model of a single component repairable system.

5 marks

Answer

Sign in to view the complete answer paper.

b

Given initial transitional rate of one component repairable model: P = [1/2 1/2; 1/4 3/4]. Evaluate the probability of being in state 2 after ‘3’ time intervals given that it is started in state 1 initially.

5 marks

Answer

Sign in to view the complete answer paper.

1010 marks
a

Explain frequency balance approach.

5 marks

Answer

Sign in to view the complete answer paper.

b

The failure rate of three components are 0.05 f/yr, 0.01 f/yr and 0.02 f/yr respectively and their average repair times are 20 hr, 15 hr and 25 hrs respectively. Evaluate the system failure rate, average repair time and unreliability if all three components must operate for system success.

5 marks

Answer

Sign in to view the complete answer paper.

OR
11

Use the frequency balance approach to evaluate the frequency of encountering and duration of residing in each of the states shown in figure 4 where rates are expressed in occ/yr.

10 marks

Answer

Sign in to view the complete answer paper.

About this paper

This page contains the question paper of RELIABILITY ENGINEERING from JNTU for B.Tech CE III III-I, November/December - 2018, R16.

For the solved question paper with answers, switch to the Answer Paper tab.

Loading Question Papers

Fetching data from database...